OpenStudy (anonymous):

Find the equation of the tangent to the graph at the indicated point. f(x) = x^2 − 2; a = 9

OpenStudy (anonymous):

Tangent is a line, so of form y=mx+c where m=f'(a) f'(x)=2x f'(9)=2*9=18 = m f(9)=9^2-2=79 = y when x=9 so we can work out c: 79=18*9+c so c=79-162=-83 ANS: y=18x-83